OpenStudy (knov):

vacuum tube belongs to one of the three parties with probabilities of 0.25: 0.5: 0.25, respectively. The probability that the lamp will work a specified number of hours, respectively equal to those parties: 0.1: 0.2: 0.4 Determine the probability that the lamp is taken at random will work a specified number of hours.

OpenStudy (knov):

@reemii

OpenStudy (reemii):

We know Part 1 : P1, P(P1) = 1/4 Part 2 : P2, P(P2) = 1/2 Part 3 : P3, P(P3) = 1/4 and P(work N hours | P1) = 0.1 P(work N hours | P2) = 0.2 P(work N hours | P3) = 0.4 The question is `P(work) = ?` It's not Bayes' formula since you don't want to compute a conditional probbility. Which is it?

OpenStudy (knov):

you mean which formula to use ?

OpenStudy (reemii):

yes

OpenStudy (knov):

there is a formula like this two people's name on it. i don't remember the name.

OpenStudy (knov):

am i close to the formula ?

OpenStudy (reemii):

The `Total probability formula` is the one to use. Were you thinking about Bienaimé-Tchebichev ?

OpenStudy (knov):

No, i think the names were different.

OpenStudy (knov):

anyways, thanks for the help again.

OpenStudy (reemii):

Np. it's "logically" the Total-probability-formula because you want to compute P(W) and you know only P(W|A), P(W|B), P(W|C), ..... where A,B,C,... divide the 'whole thing'.

OpenStudy (knov):